Quiz # 6 for Dr. Z.’s Number Theory Course for Oct. 31,, 2013

NAME: (print!) E-MAIL ADDRESS: (print!)

1. ( 4 points) Illustrate the proof of Wilson’s theorem for p = 19.

2. (3 points) How many (circular) necklaces are there of length p, with a colors? Explain!

3. (3 points) Use the Miller-Rabin primality test to investigage whether the integer 15 is prime or composite, by picking one random a’s between 2 and 13.